# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

### v1

@@ 4. Constraint Enforcement
Once users define SICs, the DPS will identify the most efficient enforcement method. The DPS initially moves all specified SICs to the earliest evaluation point to minimize unnecessary computation, akin to predicate pushdown in logical query optimization. If a SIC involves checking attributes, the DPS will move the constraint to the earliest data flow point where all attributes are defined, usually right after a semantic map or aggregate operator, since these are the only attribute-generating operators in our extended relational model. For SICs that verify operator processing logic, like soundness constraints, the DPS positions them at their associated semantic operator. Additionally, the DPS converts specified SICs into text and appends them to the LLM’s prompt, ensuring that the semantic operator’s LLM is aware of the constraints on its results. At this point, the DPS will follow either a proactive or reactive constraint enforcement strategy.
4.1. Proactive Enforcement
The DPS can proactively enforce certain types of SICs by pushing their enforcement into the LLM’s decoding process via constrained decoding (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022). In doing so, the output of the LLM will be guaranteed to satisfy the specified constraints by construction; thus, requiring no further checks on its output. Provided the constrained decoding overhead remains low, the DPS should proactively enforce SICs where feasible to optimize performance. Concerns around distorting the LLM’s distribution can be addressed with recent methods as well (Park et al., 2024).
Grammar-Based Constrained Decoding. An effective method for using constrained decoding to ensure specific SICs is by encoding them as a formal grammar. This is similar to how libraries (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) allow users to provide a JSON Schema (Org, 2010) to which the LLM’s output adheres. These libraries convert the JSON Schema into a context-free grammar or regular expression and apply a push-down or finite automaton, respectively, to guide the output. This allows the DPS to enforce certain domain constraints. Beyond exploring more efficient grammar-based decoding methods, open challenges include determining which SICs can be expressed as grammars and how to represent and compile multiple SICs into a grammar if they are defined on the same attribute.
Direct Constrained Decoding. An alternative constrained decoding method is to enforce SICs directly, without translating them into a formal grammar. Although we do not anticipate that this will allow the DPS to enforce all SICs, we expect it to enable a broader range of constraints than grammar-based methods. Previous research (Beurer-Kellner et al., 2023) introduced a direct, more flexible approach, but is significantly slower than current grammar-based techniques (Zheng et al., 2024). Whether this more flexible enforcement method can be implemented to efficiently enforce SICs is an open problem.
4.2. Reactive Enforcement and Recovery Actions
The DPS can also reactively enforce SICs by first detecting a constraint violation and then performing a recovery action.
4.2.1. Violation Detection
To detect a SIC violation, the DPS simply evaluates its associated predicate and checks if the predicate returns false. However, if multiple SICs are specified in the data flow, there are many opportunities to optimize their collective evaluation.
Constraint Reordering. If the DPS needs to execute multiple SICs sequentially (e.g., if the user specifies the constraints on the same attribute), then it will reorder the constraints to reduce execution costs. This is similar to predicate reordering in traditional database optimization. For example, the DPS can reorder SICs based on the cost and selectivity of each SIC’s predicate (Abadi et al., 2003). The cost of a SIC can be based on a single metric, such as execution time, or even a linear combination of cost metrics. The selectivity of a SIC is the probability that an output tuple fails the associated predicate. The DPS can learn the cost and selectivity of each SIC via offline or online methods by computing lightweight statistics.
Constraint Fusion. The DPS will further improve performance by merging SICs together into a single execution unit for certain scenarios. Akin to operator fusion, the DPS will combine code-based SICs together to reduce intermediate data materialization and unnecessary memory traffic. The DPS will also do the same for LLM-based SICs, where several constraints are evaluated by a single LLM invocation. Though, as more LLM-based constraints are fused together, the accuracy of results from the LLM validator is likely to degrade due to the increase in task complexity and input length (Levy et al., 2024). As such, finding the optimal number of constraints to combine that maintains target accuracy metrics while minimizing execution costs is an open problem.
4.2.2. Violation Recovery
To recover from a SIC violation, the DPS will retry the semantic operator with the failed input tuple(s) by using prompt or model-based techniques to improve the output quality. We outline a few approaches for retrying semantic operators. These approaches can be combined in different ways, and it is interesting to explore whether the DPS can learn their optimal combination with respect to different target metrics, such as accuracy, latency, and monetary cost.
Few-Shot Prompting. A common approach for retrying a semantic operator is to append the erroneous output and corresponding feedback message to the previous prompt on each failure (Singhvi et al., 2024; Shankar et al., 2024a). The feedback message can be a static error message specified for a code-based predicate or part of the response from an LLM validator. This refinement process provides the underlying LLM negative few-shot examples of the given task for in-context learning (Brown et al., 2020).
A complementary approach is to use the top- most similar positive few-shot examples, which are input-output pairs that previously satisfied all the specified SICs for the semantic operator. Importantly, examples are retrieved on the basis of their similarity to the LLM’s current input. The DPS can use a cache with a vector index to efficiently store and retrieve these examples. For each retry, the number of retrieved examples can increase to improve the quality of the LLM’s output. We believe that this approach will be particularly useful in data flow systems, where queries are mostly static. There are several challenges in our setting, e.g., determining ’s initial value and increment size for each retry, and example cache management policies.
Chain-of-Thought Prompting. When retrying a semantic operator, the DPS can also prompt the underlying LLM to use more reasoning steps in its CoT. By effectively scaling up inference-time computation, recent work has shown that long CoT in reasoning models generally improves the accuracy of responses (OpenAI, 2024a; DeepSeek-AI, 2025). For each retry, the DPS can thus prompt the LLM to use increasingly longer CoT to generate better outputs. It is unclear how long the initial CoT should be and by how much to increase it on each retry.
Model-Based Improvements. To improve the results of semantic operators, the DPS can invoke a more powerful LLM to generate a more accurate output, or use multiple LLMs to adopt a majority voting approach, e.g., where the same input is sent to increasingly more LLMs on each retry. In simple cases, such as when the operator outputs a classification label or a numerical value, the DPS can programmatically consider the outputs of each LLM and return the most common result or their average, respectively. In more complex scenarios, such as when the operator generates a summary, the DPS can use a separate LLM to consider all the outputs and produce the final result. An open avenue for exploration is to compare the accuracy of these two methods when various cost metrics are fixed.
5. Grounding Constraints
In this section, we describe grounding constraints in more detail to illustrate the design decisions required to integrate SICs. These constraints are being implemented in VectraFlow (Lu et al., 2025), an AI-augmented data flow system that we are developing at Brown. However, we aim to keep the presentation general since the ideas proposed here are applicable to other AI-augmented DPSs as well. We first discuss preliminaries related to grounding constraints and then describe their implementation in the rest of the section.
To understand the rationale for grounding constraints, we note a key distinction: an LLM’s response comes from its internal (parametric) knowledge, external (nonparametric) knowledge, or a combination of both. In an AI-augmented DPS, this external knowledge includes input tuples to the system or tuples retrieved from stored tables, termed source tuples, with their elements called source values. These values, considered factual, are essential to grounding constraints. Across extractive and abstractive contexts, users seek to confirm if a value generated by a semantic operator is anchored in source values. Grounding constraints facilitate these verifications.
5.1. Enforcement
When a grounding constraint is specified for an attribute, the DPS recursively applies checks to all ancestor attributes in its lineage. These grounding checks are necessary because the attribute may have been derived through a series of intermediate semantic operators from the source value(s). The DPS evaluates these checks immediately after attributes are generated, ensuring that the operator’s output is grounded in its input. In the example query from LABEL:lst:query, med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Since the grounding constraint specified on line 12 for med_hist_sum recursively defines a grounding check on med_hist, the query does not need to explicitly specify a grounding constraint on med_hist, though doing so is equivalent.
The notion of grounding differs depending on whether the operator is extractive or abstractive. For extractive operators, grounded output is precisely contained within the input. For abstractive operators, grounded output is factually consistent with the input. The DPS cannot inherently identify whether an operator is extractive or abstractive; hence, users must annotate extractive semantic maps. Without such annotations, attribute-generating operators in the data flow are presumed abstractive if they produce an attribute within the grounding constraint’s lineage. In the example query, we see the EXTRACTIVE keyword included for the extractive operations on lines 5 to 7, while no annotation is present for the abstractive operation that summarizes the medical history on line 8. Evaluating grounding checks on extractive and abstractive semantic operators necessitates different check types. The DPS will automatically assess the appropriate checks based on each operator’s semantics. We will now discuss these grounding check types.
5.2. Types of Grounding Checks
We propose three types of grounding checks with different verification mechanisms, targeting either extractive or abstractive tasks.
Exact Match. A grounding check based on exact match verifies whether an output attribute’s value is exactly contained within the input value. These checks are designed for extractive semantic map operations, where users require exact guarantees on output values, such as phys_exam, lab_res, and med_hist in the example query. The DPS can further optimize exact match verification by building inverted indices over stored relations. However, in data flow systems, evaluating these checks on values derived from input data streams are not amenable to such optimizations.
Factual Consistency Score. These checks evaluate whether an output value is grounded in the input value(s) by computing factual consistency scores with a specialized model. Each sentence in an output value is separately given with the input value(s) to the factual consistency model, and the resulting score determines whether the sentence is grounded. While these models are stochastic, they can validate summaries from abstractive semantic maps (e.g., med_hist_sum) and aggregates. Models like MiniCheck have also been shown to be more efficient than their LLM counterparts and reach GPT-4 accuracy (Tang et al., 2024).
LLM Validator. LLM-based grounding checks also target abstractive use cases, invoking an LLM validator to determine grounding. The LLM validator receives the output value and input value(s), and it returns a factual consistency score or a boolean value. Additionally, the LLM validator provides a feedback message, which can help the semantic operator output more grounded responses when retried. In fact, the DPS can combine the LLM-based check with the score-based check, where it invokes the LLM validator only if the factual consistency model classifies the output as ungrounded.
6. Open Problems
In this section, we present additional open problems that pave the way for promising research opportunities.
Adaptive Enforcement. A DPS can strategically manage performance and risk by dynamically applying SICs. By tracking SIC violations for each semantic operator, it determines when checks are necessary, thus reducing overhead. Violations incur a (normalized) penalty score reflecting their risk level, and these scores are combined into a sum. If this sum exceeds a set threshold, a retry is initiated to maintain critical integrity. This approach allows the system to adapt on the basis of each operator’s reliability, concentrating on critical validations while minimizing disruptions when operator confidence is high. Another approach involves adaptively enforcing SICs based on target metrics (e.g., accuracy, F1 score).
LLM Input Verification. While our discussion of SICs has mainly emphasized validating semantic operator outputs, it is equally crucial to validate the inputs, which may come from external sources (including a stored table). This step is necessary to avoid security risks, such as (i) prompt attacks (AI, 2024; Llama, 2024; Microsoft, 2024; Rebedea et al., 2023) or (ii) prevent unauthorized access to dark data (IBM, 2023), which refers to unprotected data hidden in a vector database that should not be used because it can pose security, privacy, or compliance risks due to its sensitive contents. An open challenge here is developing robust and scalable mechanisms to perform real-time, context-aware input verification that can effectively identify and filter malicious or unauthorized data before it reaches the LLM.
Enterprise-Wide Automated Constraint Verification. To enable a “check all” feature for semantic queries, each semantic operator in the query path must be automatically verified. Strategically determining what checks to use, where and how to place them involves examining data flow semantics and lineage to pinpoint where operator outputs affect subsequent results. The types of checks, which range from basic range checks to LLM validators, should match each operator’s semantics and risk of its output. Although our discussion has focused on single queries, enterprise settings often involve numerous queries that share common domain-specific constraints. Therefore, an important challenge is developing a mechanism to automatically capture and manage these recurring constraints in a constraint store and apply them consistently across all relevant semantic queries, ensuring enterprise-wide data integrity. To further enhance this process, the DPS should log all SIC violations, coupled with user feedback on incorrect validation results. This approach facilitates continuous LLM refinement through training, fine-tuning, or in-context learning, ultimately bolstering the accuracy of enterprise-wide constraint verification.
Leveraging LLM Progress. As LLM reliability and efficiency advance, they will influence SIC design and related trade-offs. Increases in LLM accuracy allow for more adaptivity in SIC enforcement. A critical question is: How do we determine when LLM accuracies warrant this shift? Additionally, as LLMs become more efficient, using advanced ensemble methods (e.g., confidence-based aggregation) to further improve reliability as well as maintaining LLM diversity to prevent shared errors become important.
@@ 7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases to ensure that modifications maintain data consistency. SICs build on top of traditional integrity constraints, improving the reliability of the semantic operators themselves by protecting against erroneous outputs from the underlying LLMs.
In ML, assertions were introduced as an abstraction to monitor and improve models, invoking UDFs over model outputs to flag possible errors (Kang et al., 2020). Recent works adopt similar ideas by further leveraging ML models and LLM validators to define guardrails on LLMs (Rebedea et al., 2023; Dong et al., 2024a; AI, 2023a, b; Index, 2023), with DSPy Assertions introducing an iterative refinement approach to improve output quality (Singhvi et al., 2024). Methods for validating an LLM’s CoT have been developed as well (Ling et al., 2023). In addition to these reactive constraint enforcement approaches, a related area of work uses constrained decoding to proactively enforce constraints (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022; Beurer-Kellner et al., 2023). SICs consolidate these various approaches into a unified and declarative abstraction designed for AI-augmented DPSs. Prior works like spade (Shankar et al., 2024b), which automatically synthesizes assertions, and EvalGen (Shankar et al., 2024c), which aligns LLM validators with human preferences, are complimentary to SICs and can be integrated to enhance usability and reliability.
@@
Several recent systems have augmented traditional database operators with the semantic capabilities of LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025; Lu et al., 2025; Shankar et al., 2024a). So far, DocETL (Shankar et al., 2024a) is the only AI-augmented DPS to provide validation as a first class citizen. They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. We further their approach by proposing a declarative, relational interface for the easy specification of common constraint classes and their optimized enforcement.
Many methods have been developed to evaluate the factual consistency of LLM responses with respect to provided source documents. This problem is also known as document-level textual entailment. In addition to an exact keyword match approach that is only appropriate for extractive use cases, recent methods for evaluating factual consistency in abstractive tasks are largely categorized as those that use LLM validators and those that use specialized factual consistency models. The latter models are specifically trained to evaluate factual consistency, primarily in the context of summarization (Tang et al., 2024; Zha et al., 2023; Kryscinski et al., 2020; Goyal and Durrett, 2020; Laban et al., 2022). These models provide factual consistency scores and are more efficient than LLM-based methods, with some even reaching GPT-4 accuracy (Tang et al., 2024; Zha et al., 2023). Regardless of the specific method for evaluating factual consistency, the interface provided by SICs is designed to be agnostic of the underlying implementation.

### v2

@@ 4. Constraint Enforcement
Constraint enforcement consists of two phases: detection and recovery. First, a constraint detects any violations in the operator’s output. Upon detection of a violation, the operator attempts to recover. This distinction aids in reasoning about optimizing constraint enforcement, as discussed in this section.
4.1. Logical Optimization
The logical optimizer takes as input a logical query plan represented as a data flow graph, where each node is either an operator or a constraint. The goal of the logical optimizer is to minimize the cost of the plan by applying rewrite rules to semantic operators and constraints. Here, we refer to cost as any arbitrary cost metric, such as latency or monetary cost. To keep the presentation focused, we only discuss logical optimizations for constraints.
Constraint Pushdown. First, the optimizer moves all constraints to their earliest evaluation point in the data flow (i.e., immediately following their corresponding operators). This allows the DPS to detect and recover from constraint violations as soon as possible rather than unnecessarily processing erroneous outputs. The optimizer also appends the textual representation of each constraint to their corresponding operator’s LLM prompt, ensuring that the underlying model is aware of the constraints on its results.
Constraint Reordering. The optimizer then reorders the SICs for the same operator. Following prior work (Abadi et al., 2003), the reordering is based on the cost and selectivity of each constraint, where selectivity is defined as the probability that a tuple satisfies the constraint’s predicate. Cost and selectivity estimates can be computed from sampled data. We discuss more details on estimation in Section 4.2.2.
4.2. Physical Optimization
Given an optimized logical plan, the physical optimizer considers various implementations of the plan and returns the optimal one. We begin with a discussion of different approaches to constraint enforcement, and then outline ideas for optimized plan selection.
4.2.1. Enforcement Implementations
Constraints can be enforced in multiple ways, typically categorized as reactive or proactive, based on whether enforcement occurs after or during response generation, respectively. While the DPS can enforce all constraints reactively, only certain ones can be enforced proactively.
Reactive Enforcement. After the LLM produces its entire response, there are different methods to detect a constraint violation. For example, a grounding constraint on an extractive operator can simply check if the operator’s output attribute is contained in the input. On the other hand, a grounding constraint on an abstractive operator can invoke a single LLM judge (Zheng et al., 2023), a fact-checking model (Tang et al., 2024), a compound AI system (Chern et al., 2023), or anything in between. Similarly, there are different approaches for an operator to recover from a violation by retrying. The common approach used in current systems is iterative refinement with feedback (Singhvi et al., 2024; Shankar et al., 2024a; Madaan et al., 2023). Furthermore, the operator can retrieve additional semantically similar few-shot examples (Brown et al., 2020) of prior input-output pairs that satisfied the constraint. These examples can be stored in a cache built on a vector index. Other approaches include retrying with more powerful models, larger ensemble sizes, or higher thinking budgets (Anthropic, 2025; Google, 2025).
Proactive Enforcement. Proactively enforcing constraints can lower costs significantly. For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, identifying and masking invalid tokens to ensure compliance by construction. Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token validation. However, constrained decoding is not limited to grammar-based approaches. For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints. Yet, open source libraries currently do not offer this functionality.
An open challenge is to identify which other SICs can be enforced via constrained decoding. Moreover, developing a framework to efficiently apply various constraints on different output attributes also remains unresolved. Prior research (Beurer-Kellner et al., 2023) has seen some advancement but remains slower than modern grammar-based techniques (Zheng et al., 2024). Nonetheless, it provides a foundation for future exploration. Importantly, we suggest that the DPS should limit constrained decoding to simple constraints (seen by the LLM during training) to avoid sampling unlikely token distributions. This rationale also underlies why constraints are included in the LLM’s prompt before decoding.
Besides deterministic constrained decoding, the DPS can proactively enforce stochastic constraints by asynchronously detecting violations in partial streamed outputs. This approach is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint. For instance, the DPS can proactively enforce a grounding constraint on an abstractive operator by asynchronously validating generated summary sentences with a fact-checking model (Tang et al., 2024). Upon a violation, the operator can backtrack to the decoding step right before the violation. Alternative techniques like re-prompting with partial feedback are also worth considering to address this problem.
4.2.2. Optimized Plan Selection
For a fixed set of constraints and their implementations, the physical optimizer can adapt existing techniques (Liu et al., 2025b; Russo et al., 2025) to select an optimal query plan defined by user-specified cost or reliability (a.k.a. quality) thresholds (omitted from the query interface for brevity). Specifically, rather than measuring the reliability of an operator’s implementation in terms of validation data or the best implementation from prior beliefs (e.g., a “champion model”), reliability is instead defined by the proportion of output tuples that satisfy all of the operator’s constraints (after necessary retries). However, the challenge is selecting each constraint’s implementation to begin with.
As discussed in Section 4.2.1, there are many ways to detect a constraint violation, each with different cost-reliability trade-offs. Formally, the cost of a constraint’s implementation is the overhead of detecting a violation in a single output. The reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation. A deterministic constraint is completely reliable (i.e., 100% precision and recall), while a stochastic constraint is not completely reliable, due to invoking models such as LLMs.
To choose the best constraint implementation, we must estimate costs and reliability for each option. While cost estimation is often simple and can be done during runtime with sampled data, it might not always be necessary. For instance, proactively enforcing a deterministic constraint through constrained decoding is usually quicker than doing it reactively.
Assessing the reliability of stochastic constraints is more challenging. Unlike deterministic constraints that are always reliable, the reliability of stochastic constraints depends on (1) how they are implemented, (2) the output tuples from upstream operators in the query plan, and (3) the input tuples to the query. Moreover, obtaining robust reliability estimates for constraint implementations can be difficult, as they often need ground-truth labels from human annotators. We present several strategies to address this issue.
One strategy is to use historical data and workloads to estimate the precision and recall of different implementations for each constraint, providing these estimates to the physical optimizer for use at runtime. This approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates. Since metrics are averaged across diverse query contexts, they do not effectively reflect interactions with upstream operators. Furthermore, data and workload drift may make these estimates obsolete. It is also generally costly for organizations to annotate execution traces.
An alternative approach is to develop a shared foundation model for organizations, where given representations of (1) a constraint’s implementation, (2) the implementation of the query plan up until the constraint, and (3) the underlying data, the model predicts the precision and recall for the constraint’s implementation. Organizations can optionally fine-tune the model for their use cases. Recent work on foundation database models (Wehrstein et al., 2025) are applicable here, with extensions to address the complexities of semantic processing.
A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., expected confidence) of a constraint’s implementation at runtime based on sampled data. Rather than defining precision and recall thresholds, users specify confidence thresholds instead. Although many conventional classifiers inherently provide confidence scores, they can be derived from LLMs by applying softmax to the “True” and “False” token logits. The difficulty lies in achieving calibrated confidence scores that accurately represent the implementation’s true correctness probability. Studies (Kadavath et al., 2022; Liu et al., 2025a) indicate that methods like temperature scaling (Guo et al., 2017) can enhance calibration in binary classification tasks. Though, further research is needed to obtain highly-calibrated constraints.
5. Observability
After a user submits a query to the DPS, the observability stack enables the user to understand the performance of their query’s operators and constraints. The stack has three components. The observability interface allows the user to view relevant metrics (e.g., cost, reliability) and inspect query results. The observability store maintains metrics and outputs from each stage in the query. Lastly, the labeling interface enables human annotators to provide ground-truth labels to compute the precision and recall of each stochastic constraint. While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread in both academia and industry (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), their reliability remains uncertain. As a result, labeling interfaces remain essential. A key open challenge lies in ensuring that the observability store can support the proposed interfaces efficiently at scale and that these interfaces are effectively designed. We detail each component of the stack below.
Observability Interface. Given a user-specified query ID, the observability interface immediately shows metrics such as cost, operator reliability, and constraint selectivity. However, to observe the reliability for stochastic constraints, the user must submit a labeling request for their query. Once the labeling is complete, the interface displays the precision and recall of each constraint. To avoid confusion, the interface indicates whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well. The observability interface also allows the user to inspect individual output tuples. A tuple is annotated if any upstream operator violated a constraint but CONTINUEed to propagate the error. The user can view the tuple’s lineage for more fine-grained debugging.
Observability Store. The observability store is modeled as additional relations in the DPS. It maintains the lineage of each query, enabling systematic debugging and performance monitoring. To obtain human annotations, the DPS can draw concepts from crowd-sourced databases (Franklin et al., 2011; Marcus et al., 2011; Park et al., 2012), such as crowd-sourced attributes and user interface generation; however, the DPS need not implement these ideas in their full generality. For example, suppose the DPS stores a relation where each tuple represents an invocation of a constraint. The relation includes attributes for the constraint’s input and output (i.e., predicted label) among others (e.g., query ID). In particular, the relation also has a crowd-sourced attribute for the true label (initially null). When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample of the relation) the precision and recall of each constraint for the given query ID. The query populates each sampled true label by sending tasks to a crowd-sourcing platform.
Labeling Interface. A key aspect of crowd-sourced databases is their ability to automatically create user interfaces for crowd workers based on declarative queries. Here, the DPS produces a labeling interface for annotators, which displays a description of the constraint and its inputs. The annotator’s role is to determine if the inputs meet the constraint. To avoid bias, the original predicted label should be hidden.
6. Enterprise-Wide Constraints
The introduction of SICs into an enterprise environment presents both new opportunities and challenges. To effectively manage SICs within an organization, we propose the development of a constraint store and outline its key desired features.
Auto-Guardrails and Recommendations. Although declaring SICs is relatively straightforward, determining the applicable constraints is more difficult. This requires anticipating the potential failure cases for each semantic operator in the query. Recent systems have been developed to assist users in formulating constraints (Shankar et al., 2024b, c), but these approaches typically generate suggestions based solely on the contents of the user’s query. In enterprise settings, constraint recommendations can be greatly enriched. Organizations typically enforce domain-specific rules that necessitate certain constraints for all pertinent queries. For instance, a hospital might require excluding PII in queries unless essential for analysis. Moreover, different teams within the same organization might use similar queries with shared constraints. Thus, the constraint store needs to automatically capture and manage common constraints organization-wide, facilitating their reuse and recommendation in related queries. We suggest using text embeddings (OpenAI, 2022; Reimers and Gurevych, 2019) and graph embeddings (Xu et al., 2019; Kipf and Welling, 2017; Veličković et al., 2018) to represent constraints and queries. This strategy boosts the efficiency of similarity searches (Malkov and Yashunin, 2020) for identifying relevant constraints, which can then be refined by an LLM to match the user’s context.
Conflicts. Conflicts can arise when queries contain multiple SICs (e.g., from accumulating shared constraints across an organization’s workloads). For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy both constraints without redaction or abstraction. Likewise, domain constraints over disjoint value sets can lead to inevitable failures. The constraint store allows for conflict analysis using static reasoning and LLM-based methods, issuing warnings. Users can address conflicts by adjusting priorities (e.g., making one constraint as soft), relaxing scope, or merging constraints as needed. This approach builds on early research in active databases (Widom, 1996; Hanson, 1996; Dayal et al., 1988), which suggested techniques for detecting rule conflicts and execution issues, such as non-termination, before execution (Aiken et al., 1992).
@@ 7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases.
ML assertions were introduced to monitor and enhance model performance by applying UDFs to model outputs to find potential errors (Kang et al., 2020). Building upon this idea, recent developments broaden assertions for LLMs by employing LLMs as judges and ML classifiers for both offline evaluation (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024). Tools like spade (Shankar et al., 2024b) and EvalGen (Shankar et al., 2024c) guide users by suggesting relevant assertions and their implementations (e.g., Python functions, LLM grader prompts). These tools complement SICs and can suggest declarative constraints to boost enforcement reliability and efficiency. Our vision for enterprise constraints also draws inspiration from these efforts. DSPy Assertions enable users to define assertions in LLM pipelines using UDFs, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024). While SICs use iterative refinement, they do not currently include prompt optimization capabilities, although they can be added. To the best of our knowledge, no framework provides a fully declarative method for optimized constraint enforcement.
A related area of work is constrained decoding, which proactively enforces constraints so that the LLM’s response adheres to them by construction (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). Current frameworks focus on grammar-based approaches; however, similar to earlier work (Beurer-Kellner et al., 2023), we view constrained decoding as a more general approach that is not limited to expressing constraints as grammars.
@@
Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025). So far, DocETL is the only AI-augmented DPS to provide validation as a first class citizen (Shankar et al., 2024a). They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. However, their approach to validation is largely imperative. Palimpzest does not support constraints, but they allow users to specify quality thresholds to select optimized physical plans (Liu et al., 2025b; Russo et al., 2025).